# Data Mesh: a Systematic Gray Literature Review

### v1

@@ 1. Introduction
The world is living in the golden age of data. IDC (International Data Corporation) predicts that by 2025 the amount of digital data generated by enterprises and individuals will grow 61% to 175 zettabytes (Rydning et al., 2018). While the advances in the analytic data landscape can enable organizations to turn their ever-growing raw data into value, the difficulties in the integration, management, and governance of data at scale hamper the success of data analytic projects of organizations (Ramachandran, 2020; Janssen et al., 2020a). In the current practices of centralized data management, domain teams rely on a central data management team to collect, process, and manage domain data. This central team is becoming the bottleneck for unlocking the value from domain data. There is a need to shift the domain data’s responsibility from the central team to domain teams (Dehghani, 2022).
@@
A data mesh is emerging as a novel decentralized approach for managing data at scale by applying domain-oriented, self-serve design and product thinking (Dehghani, 2022). Zhamak Dehghani first defined the term data mesh in 2019 ((2019), Thoughtworks). Figure 1 shows the search index for “Data Mesh” on Google Trends for the past five years. A clear increasing trend line can be seen, indicating the increased interest in data mesh. Despite the popularity of data mesh, little academic literature exists on the topic. However, as organizations increasingly look into data mesh, we saw an opportunity to systematically review the gray literature to characterize the concept better and drive a research agenda.
Figure 1. Google trends for the word “Data Mesh”.
@@
Zhamak (Dehghani, 2022) defines the architecture of data mesh based on four principles; data as products, domain ownership of data, self-serve data platform, and federated computational governance. The data assets in an organization are offered as products that enable consumers to efficiently and autonomously discover and use domain data. Domain teams closest to the origin of data create and manage data products. A mesh of data products emerges when domains consume data products from each other. To ensure the data products are interoperable and compliant with internal and external regulations and policies, a federated governance team is responsible for defining and enforcing product interoperability standards and compliance policies across the data mesh. Finally, a self-serve data platform supports the domains by offering infrastructure and platform services necessary to build, maintain and manage data products.
@@
To a certain extent, we observed a strong analogy between data mesh and service-oriented architecture (SOA), particularly with domain-driven microservices (Dehghani, 2022; (2019), Thoughtworks; Christ et al., 2022). The success of the domain-oriented design and microservices architecture was an inspiration for inventing the data mesh architecture ((2019), Thoughtworks; Dehghani, 2022). Data products are often exposed as services and consumed through APIs. A data mesh relates and connects data products, which is analogous to a service mesh. The product owners can publish their data products, and consumers can discover and interconnect the desired products to perform cross-domain data analysis and build new value-added data products, which is analogous to service composition.
@@
The remainder of this article is organized as follows. Section 2 describes our systematic literature review methodology. Based on the findings of the gray literature, Section 3 defines the data mesh and its four principles and presents its potential benefits and drawbacks. Section 5 describes our three reference architectures for the data mesh, while Section 6 lists potential research challenges. Section 7 summarizes the academic research studies on data mesh and the literature review studies relevant to our survey. Finally, we discuss the threats to the validity of our research in Section 8 and conclude the paper in Section 9.
@@ 2. Methodology
Due to the novelty of the data mesh paradigm, little academic literature is available on the topic. At the same time, since the data mesh concept has its roots in the industry as opposed to academics ((2019), Thoughtworks), gray literature about the topic, such as blog posts and white papers, is widely available. Since the quality of structured literature reviews (SLRs) depends greatly on the availability of enough source material, a gray literature review (GLR) was conducted to investigate the state-of-the-art for data mesh. In this work, we follow the guidelines provided by Garousi et. al. (Garousi et al., 2019), who adapted the SLR guidelines of Kitchenham and Charters (Keele et al., 2007) specifically for GLRs. Following these methodologies, our GLR is organized into three stages: a planning and design phase, an execution phase, and a reporting phase. A high-level overview of these phases and their steps are depicted in Figure 2.
@@ 2.1. Planning and Design of the Gray Literature Review
Our main goals were to provide a unified definition of the data mesh concept, create reference architectures for data mesh, and identify research challenges. We formulated four research questions to achieve these goals:
RQ1. What is a data mesh and why is it needed?:
A data mesh embodies four design principles ((2019), Thoughtworks; Dehghani, 2022): data as a product, domain-driven ownership of data, self-serve data platform, and federated computational governance. Practitioners may have different perspectives on these principles and experiences applying them. Hence, RQ1 aims to understand and characterize the four principles of data mesh from a practitioner’s standpoint.
RQ2. What are the benefits and concerns of adopting data mesh?:
A data mesh has advantages and disadvantages and is not a one-size-fits-all approach for data management. Thus, RQ2 aims to understand data mesh’s potential benefits, implementation challenges, and drawbacks.
@@
Some organizations have started their data mesh journey and report their experience, including development and operation guidelines, options, and examples. Hence, RQ3 aims to analyze the relevant gray articles to create a set of reference architectures to guide organizations in making better design choices embarking on their data mesh journey. Furthermore, due to the novelty of the data mesh architectures, we aim to build reference architectures by mapping findings from the gray literature into the architectures used in comparable domains.
@@
In the gray articles, practitioners report critical challenges; hence, RQ4 aims to identify research challenges concerning data mesh based on practitioners’ challenges.
Study Selection Strategy.
@@
Initial Search.
As is common practice in GLRs, our search string selection process was based on trial queries with different search strings, which can indicate the usefulness of that search string (Straus and Corbin, 2014). After trying several combinations of search terms, we settled on two queries that appeared to yield promising results for answering our research questions:
Q1::
@@
Q2::
Decentralized Data Architecture
These queries were then executed in the Google search engine111google.com on September $1^{st}$ 2021, and all search results were checked against our inclusion and exclusion criteria. Due to the relative novelty of data mesh, information and sources at that time were mostly limited to the generic data mesh principles. Therefore, the GLR was extended to include data sources about the architectural design of data mesh until May $1^{st}$ 2022. This step resulted in 155 data sources from 2019 until 2022.
Inclusion and Exclusion Criteria.
@@
Next, works that consist primarily of subjective statements that promote the author’s opinion, as opposed to factual information, were excluded. Furthermore, we checked whether the publishing organization is respected in the field of information technology and whether the author is a reputable individual, i.e., they are associated with a reputable organization, they have reputable expertise, or have published other articles in the field of data management.
@@
The final selection consists of 114. Figure 4 shows the distribution of the selected sources. A full overview of each source can be found in our online appendix (see Section 2.4).
Figure 4. The publication date of the selected sources (114). The significant upward trend in publications indicates the novelty of data mesh.
2.2. Conducting the Gray Literature Review
@@
The second set of steps is related to the execution of the review in line with a qualitative analysis methodology. Structural coding was used to get an initial overview of the topics, after which descriptive coding was used to summarize parts of the topics. Structural coding was used to identify larger text segments on a broad topic. Descriptive coding was then used to go more in-depth on these topics by summarizing the basic topic into a word, or short phrase (Saldaña, 2021). Coding was performed using Atlas.ti too  [^2] https://atlasti.com/.
@@
Like (Kumara et al., 2021), an initial set of 20 sources was randomly selected and analyzed to establish an initial set of codes. These were discussed between the first three authors of the paper. Once consensus on the initial set of codes was reached, the entire dataset was coded. The grouping of codes into topics and the identification of the categories were established through discussion by the first three authors of this paper. An example of codes for a single topic of the “Domain Ownership” category can be seen in Table 1
@@
We first used the data extracted from the gray literature using the established codes to answer RQ1 and RQ2. Next, to address RQ3, we created three reference architectures by mapping our findings to the reference architectures used in service-oriented computing (Papazoglou, 2008; Papazoglou et al., 2007; Li et al., 2019; Zhou et al., 2002; Kumara et al., 2017). The first three authors of the paper collaboratively created the reference architectures over a set of meetings. The other authors validated such architectures. We addressed any discrepancies through discussions. We followed a similar approach to answer RQ4. We identified the potential research challenges by mapping the difficulties mentioned by the practitioners in the gray articles to the relevant research issues from the service-oriented computing (Papazoglou et al., 2007; Wei and Blake, 2010; Papazoglou and Van Den Heuvel, 2007) and data management (Dustdar et al., 2012; Stonebraker et al., 2018; Al-Ruithe et al., 2019; Janssen et al., 2020a).
2.3. Reporting the Results
@@
The results from the systematic gray literature are presented in the rest of this paper. We defined four main categories as shown in Figure 5. For each category, a table shows the topics, the underlying atomic codes, and the sources that address these codes. Table 1 shows a snippet of the results table for the Domain Ownership category. Each topic has several underlying codes. For each topic and code, we provide a unique ID, the frequency of their occurrence in the sources, and the sources themselves. Note that a particular code can appear multiple times in a single source.
Table 1. A snippet of the results for the “Domain Ownership” category.
Sub Category Atomic Codes
ID Name Frequency ID Name Frequency Unique Sources
2.1 Domain Formation 42 2.1.1 Decentralized Domains 14 S1,S9,S19,S21,S22,S42,S45,S48,S52,S55,S78,S91,S110
2.1.2 Domain-Driven Design Method 28
S2,S11,S36,S37,S39,S45,S46,S53,S55,S58,S71,S75,S78,
S79,S87,S90,S91,S96,S98,S100,S103,S110,S111,S112
Figure 5. Topics discussed in the analyzed studies.
@@ 2.4. Replication Package
We made the related data available online [^3] https://github.com/IndikaKuma/DataMesh to enable further validation and replication of our study. It contains the full list of sources and the qualitative analysis (i.e., codes, groupings, and analysis) performed with the Atlas.ti tool [^4] https://atlasti.com/.
3. RQ1. What is Data Mesh and Why is it Needed?
@@
In this section, based on the results from our GRL, we first define a data mesh, explaining its four principles: (i) data product thinking (Section 3.1), (ii) domain ownership (Section 3.2), (iii) federated computational governance (Section 3.3), (iv) and self-serve data platform (Section 3.4). Next, we explain the expected benefits of a decentralized data mesh architecture, the concerns of a data mesh implementation, and to which organizations a data mesh architecture is applicable (Section 4).
Tables 2, 3, 4, 5 and 6 overviews the sub-categories derived from the gray literature review. Each table also overviews the atomic codes, their frequency, and the unique studies that mentioned them. The results refer to the sub-categories or codes through their labels; for example, ID1.1.1 refers to the code “Data” belonging to the sub-category “Components” part of the category “Data As a Product”.
Table 2. Data as a product: codes and sources.
Sub Category Atomic Codes
ID Name Frequency ID Name Frequency Unique Sources
1.1 Components 67 1.1.1 Data 13 S1, S24, S37, S57, S82, S97-S99, S103, S104, S108, S110
1.1.2 Metadata 13 S1, S10, S37, S45, S53, S57, S82, S97, S98, S110, S113
1.1.3 Code 13 S1, S37, S86, S97, S98, S102, S103, S104, S108, S110
1.1.4 Interfaces 19 S1,S24,S28,S29,S37,S53,S67,S82,S86,S88,S103,S108,S110,S113
1.1.5 Infrastructure 9 S1,S37, S53, S57, S77, S98, S97
1.2
Data Product
Types
32 1.2.1 Basic Data Products 12 S2,S6,S20,S86-S88,S96,S106,S109,S110,S112
1.2.2 Composite Data Products 20 S2,S6,S86-S88,S95,S96,S98,S106,S109,S110,S112
1.3
Characteristics
of
Data Products
132 1.3.1 Discoverable 28
S1,S2,S3,S7,S9,S15,S16,S32,S33,S37,S38,S43,S53,S55,
S57,S65,S73,S75,S77,S82,S88,S99,S100
1.3.2 Interoperable 17
S2,S7,S15,S32,S37,S47,S53,S57,
S63,S65,S77,S88,S99,S100,S113
1.3.3 Natively Accessible 12 S15,S16,S32,S33,S37,S53,S66,S82,S88,S91,S97,S99
1.3.4 Self-describing 12 S2,S3,S7,S28,S37,S38,S53,S57,S65,S77,S88,S100
1.3.5 Understandable 9 S1,S9,S15,S16,S19,S55,S66
1.3.6 Secure 21 S1,S7,S9,S15,S16,S18,S19,S24,S32,S37,S53,S57,S65,S77,S82
1.3.7 Trustworthy 22
S1,S2,S7,S9,S15,S16,S32,S37,S38,S53,S57,S65,S77,S82,
S88,S97,S99,S100,S113
1.3.8 Valuable 11 S15,S16,S85,S91,S109-S112
3.1. Data as a Product
@@
This principle applies product thinking to analytical data to offer the data as a valuable asset to potential data consumers.
ID1.1. Components.
Similar to a service in SOA, a data product is an autonomous entry that receives or takes the input data from one or more sources, applies some computation/transformation on the received data, and produces and serves the results excepted by the consumers. We identified five common elements for a data product: data, metadata, code, interfaces, and infrastructure. However, some data products may include product-specific artifacts, e.g., a machine-learning model.
The metadata of a data product may describe its owners, schemas, quality metrics, and access policies. The product code is responsible for ingesting/consuming and transforming the data from source systems or upstream data products and serving the transformed data to end-users or downstream data products. The product can also include code to enforce various data governance policies programmatically (policy as a code), including data quality, access control, and privacy. A product also has various interfaces that define and enable interactions with other data products, data sources, data sinks, and management applications. Lastly, a product needs infrastructure resources and software platform services to ingest and store its data, deploy and execute its code, and make its data and capabilities accessible through its interfaces.
ID1.2. Types of Data Products.
@@
We identified two types of data products: atomic data products and composite data products. The first type uses data from data source systems (e.g., operational databases, external APIs, and sensors) as its source data and transforms it into the desired format. The processed data is then provisioned as a product to be consumed by the organization’s end-users or downstream data products. Gray literature refers to atomic data products as source-aligned because they mostly ingest data from the operational source systems. Composite data products use other data products (both basic or composite data products) for their source data and integrate them to create new value-added products. Gray literature also categorizes composite data products as aggregates and consumer-aligned data products. Compared to the aggregate, customer-aligned data products are generally not sold as data products to other domains; instead, they serve data to implement the end-user use cases of the organization readily.
@@
Figure 6. An example of data products in different domains, adopted from (Christ et al., 2022).
@@
Figure 6 shows data products spread across domains. Some data products, e.g., product search, order, and invoice, are atomic products. They mostly ingest raw data from the underlying operational data sources (e.g., product and order databases), clean and transform the raw data, and store them locally. They also provide the interfaces to serve their local domain data to upstream data products. The composite data products, e.g., 360-degree customer views, product recommendations, and funnel analytic dashboards, aggregate data from atomic or other composite data products. Some data products, e.g., funnel analytics and CRM dashboards, primarily serve end-users.
ID1.3. Characteristics of Data Products.
From the gray literature, we identified eight attributes of a high-quality data product.
ID1.3.1. Discoverable.:
Discoverability aims to overcome the high costs and friction of data discovery in centralized data architectures. Data consumers should be able to discover the data products in an organization easily. Gray literature recommends using a central data/product catalog. All products and their metadata should be published in the catalog so consumers can browse and search for data products easily.
ID1.3.2. Interoperable.:

### v2

@@ 1. Introduction
The world is living in the golden age of data. IDC (International Data Corporation) predicts that by 2025, the amount of digital data generated by enterprises and individuals will grow 61% to 175 zettabytes (Rydning et al., 2018). While the advances in the analytic data landscape can enable organizations to turn their ever-growing raw data into value, the difficulties in the integration, management, and governance of data at scale hamper the success of data analytic projects of organizations (Ramachandran, 2020; Janssen et al., 2020). In the current practices of centralized data management, domain teams rely on a central data management team to collect, process, and manage domain data. This central team is becoming the bottleneck for unlocking the value from domain data. There is a need to shift the domain data’s responsibility from the central team to domain teams (Dehghani, 2022a).
@@
A data mesh is emerging as a novel decentralized approach for managing data at scale by applying domain-oriented, self-serve design and product thinking (Dehghani, 2022a). Zhamak Dehghani first defined the term data mesh in 2019 (Dehghani, 2019b). Figure 1 shows the search index for “Data Mesh” on Google Trends from January 2019 to December 2023 (monthly average). The interest in data mesh grew gradually until early 2022. Then, there was a sharp increase, followed by a generally sustained popularity. Despite its popularity, little academic literature exists on the data mesh topic. However, as organizations increasingly investigate data mesh, we saw an opportunity to systematically review the gray literature to better characterize the concept and drive a research agenda.
[…]
Google trends for the term “Data Mesh”.
Figure 1. Google trends for the term “Data Mesh”.
@@
Zhamak (Dehghani, 2022a) defines the architecture of data mesh based on four principles: data as a product, domain ownership of data, self-serve data platform, and federated computational governance. The data assets in an organization are offered as products that enable consumers to efficiently and autonomously discover and use domain data. Domain teams closest to the origin of data create and manage data products. A mesh of data products emerges when domains consume data products from each other. To ensure the data products are interoperable and compliant with internal and external regulations and policies, a federated governance team is responsible for defining and enforcing product interoperability standards and compliance policies across the data mesh. Finally, a self-serve data platform supports the domains by offering infrastructure and platform services necessary to build, maintain, and manage data products.
@@
To a certain extent, we observed a strong analogy between data mesh and service-oriented architecture (SOA), particularly with domain-driven microservices (Dehghani, 2022a, 2019b; Jochen Christ and Harrer, 2022). SOA represents a software application as a set of loosely coupled collaborating services (Papazoglou and Van Den Heuvel, 2007). It promotes the idea of building applications by composing existing applications from multiple parties through discovering and invoking network-available services. The microservices architecture is a variant of SOA. The success of the domain-oriented design and microservices architecture was an inspiration for inventing the data mesh architecture (Dehghani, 2022b, 2019b, a). Data products are often exposed as services and consumed through APIs. A data mesh relates and connects data products, which is analogous to a service mesh. The product owners can publish their data products, and consumers can discover and interconnect the desired products to perform cross-domain data analysis and build new value-added data products, which is analogous to service composition.
@@
The remainder of this article is organized as follows. Section 2 describes our systematic literature review methodology. Based on the findings of the gray literature, Section 3 explores the data mesh and its four principles and Section 4 presents its potential benefits and drawbacks. Section 5 describes our three reference architectures for the data mesh, while Section 6 lists potential research challenges. Section 7 summarizes the academic research studies on data mesh and the literature review studies relevant to our survey. Finally, we discuss the threats to the validity of our research in Section 8 and conclude the paper in Section 9.
@@ 2. Methodology
Due to the novelty of the data mesh paradigm, little academic literature is available on the topic. At the same time, since the data mesh concept has its roots in the industry as opposed to academics (Dehghani, 2022b, 2019b), gray literature about the topic, such as blog posts and white papers, is widely available. Since the quality of structured literature reviews (SLRs) depends greatly on the availability of enough source material, a gray literature review (GLR) was conducted to investigate the state-of-the-art for the data mesh. In this work, we follow the guidelines provided by Garousi et. al. (Garousi et al., 2019), who adapted the SLR guidelines of Kitchenham and Charters (Keele et al., 2007) specifically for GLRs. Following these methodologies, our GLR is organized into three stages: a planning and design phase, an execution phase, and a reporting phase. A high-level overview of these phases and their steps are depicted in Figure 2.
[…]
Systematic gray literature review process.
@@ 2.1. Planning and Design of the Gray Literature Review
Our main goals were to provide a unified definition of the data mesh concept, create reference architectures for the data mesh, and identify research challenges. We formulated four research questions to achieve these goals:
RQ1. What is a data mesh?:
A data mesh embodies four design principles (Dehghani, 2022b, 2019b, a): data as a product, domain-driven data ownership, self-serve data platform, and federated computational governance. Practitioners may have different perspectives on these principles and experiences applying them. Hence, RQ1 aims to understand and characterize the four principles of data mesh from a practitioner’s standpoint.
RQ2. Why is the data mesh needed?:
Specifically, what are the benefits and concerns of adopting it, and when should it be used? A data mesh has advantages and disadvantages and is not a one-size-fits-all approach for data management. Thus, RQ2 aims to understand the data management issues that data mesh tries to address and the potential benefits, use cases, implementation challenges, and drawbacks of data mesh.
@@
Some organizations have started their data mesh journey and report their experience, including development and operation guidelines, options, and examples. Hence, RQ3 aims to analyze the relevant gray articles to create a set of reference architectures to guide organizations in making better design choices when embarking on their data mesh journey. Furthermore, due to the novelty of the data mesh architectures, we aim to build reference architectures by mapping findings from the gray literature into the architectures used in comparable domains.
@@
Practitioners report critical challenges in the gray articles; hence, RQ4 aims to identify research challenges concerning data mesh based on practitioners’ challenges.
[…]
The search and selection process. Steps are shown sequentially, and each step reduces the number of candidate sources until the final selection. After the final exclusion step, an inter-rater reliability test—featuring the well-established Cohen Kappa coefficient calculation—was performed with $\kappa=0.79$ to ensure that reviewer bias was not inappropriately high.
@@
As is common practice in GLRs, our search string selection process was based on trial queries with different search strings, which can indicate the usefulness of that search string (Straus and Corbin, 2014). We started with the keywords used in our research questions. We kept improving the combination until we settled on the following query, which yielded promising results for answering our research questions.
Query::
@@
The query was then executed in the Google search engine on September $1^{st}$ 2021, and all search results were checked against our inclusion and exclusion criteria. Due to the relative novelty of data mesh, information and sources at that time were mostly limited to the generic data mesh principles. Therefore, the GLR was extended to include data sources about the architectural design of data mesh until May $1^{st}$ 2022. This step resulted in 155 data sources from 2019 until 2022.
@@
Next, works that consist primarily of subjective statements that promote the author’s opinion, as opposed to factual information, were excluded. Furthermore, we checked whether the publishing organization is respected in information technology and whether the author is a reputable individual, i.e., they are associated with a reputable organization, have reputable expertise, or have published other articles in the field of data management.
@@
The final selection consists of 114. Figure 4(a) shows the distribution of the selected sources. A full overview of each source can be found in our online appendix (see Section 2.3).
(a) The publication date of the selected sources.
(b) Topics discussed in the analyzed studies.
Figure 4. Source statistics.
[…]
Source statistics
2.2. Conducting and Reporting the Gray Literature Review
@@
The second set of steps is related to the execution of the review in line with a qualitative analysis methodology. Structural coding was used to get an initial overview of the topics, after which descriptive coding was used to summarize parts of the topics. Structural coding was used to identify larger text segments on a broad topic. Descriptive coding was then used to go more in-depth on these topics by summarizing the basic topic into a word or short phrase (Saldaña, 2021). We leveraged Atlas.ti [^1] https://atlasti.com/ for the coding process.
@@
Like (Kumara et al., 2021), an initial set of 20 sources was randomly selected and analyzed to establish an initial set of codes. These were discussed between the first three authors of the paper. The entire dataset was coded once consensus on the initial set of codes was reached. The grouping of codes into topics and the identification of categories were established through discussion by the first three authors of this paper.
@@
We first used the data extracted from the gray literature using the established codes to answer RQ1 and RQ2. Next, to address RQ3, we created three reference architectures by mapping our findings to the reference architectures used in service-oriented computing (Papazoglou, 2008; Papazoglou et al., 2007; Li et al., 2019; Kumara et al., 2017). The first three authors of the paper collaboratively created the reference architectures over a set of meetings. The other authors validated such architectures. We addressed any discrepancies through discussions. We followed a similar approach to answer RQ4. We identified the potential research challenges by mapping the difficulties mentioned by the practitioners in the gray articles to the relevant research issues from the service-oriented computing (Papazoglou et al., 2007; Wei and Blake, 2010; Papazoglou and Van Den Heuvel, 2007) and data management (Dustdar et al., 2012; Stonebraker et al., 2018; Al-Ruithe et al., 2019; Janssen et al., 2020).
@@
The results from the systematic gray literature are presented in the rest of this paper. We defined four main categories as shown in Figure 4(b). In the online appendix, we provide the details of each category, including topics, underlying atomic codes, and the sources that address these codes.
@@ 2.3. Replication Package
We made the related data available online [^2] https://github.com/IndikaKuma/DataMesh to enable further validation and replication of our study. It contains the full list of sources, the qualitative analysis (i.e., codes, groupings, and analysis) performed with the Atlas.ti tool, and code tables.
3. RQ1. What is Data Mesh
@@
Zhamak (Dehghani, 2022b, 2019b, a) defines the data mesh concept using four principles: (i) data product thinking, (ii) domain ownership of data, (iii) federated computational governance, (iv) and self-serve data platform. Based on the findings from our gray literature, Sections 3.1, 3.2, 3.3 and 3.4 aim to study the practitioners’ perspective on these principles and how they expand or implement the principles.
3.1. Principle 1: Data as a Product
@@
This principle applies product thinking to analytical data, offering it as a valuable asset to potential data consumers.
Elements of a Data Product.
Zhamak (Dehghani, 2022b, 2019b, a) defines a data product as an architectural quantum (an independently deployable, high-cohesive component encompassing all the structural elements required for its function (Ford et al., 2022)). A data product is a composition of all its key elements: 1) data and metadata, 2) code for ingesting, processing, serving, and governing data, and 3) infrastructure for building, deploying, and executing code and storing data and metadata. The metadata of a data product may include information such as product owners, access endpoints, data models, access policies, data quality metrics, and data lineage. A data product is a node in the data mesh and an autonomous entry that receives or takes the input data from one or more sources, applies some computation/transformation on the received data, and produces and serves the results expected by the consumers.
While most gray literature generally agrees with Zhamak’s definition of a data product as an architectural quantum, some sources (Nivas Shankar and Hasson, 2021; Hicks, 2021) consider a curated data set (without code and infrastructure) a data product. Several sources also introduced the concept of a data contract (Strengholt, 2021; Bulpin, 2021; Raab and Chiapusso, 2021; Jochen Christ and Harrer, 2022; Wolf, 2022; Strengholt, 2022a) to represent the interface of a data product. It can include data access endpoints and the syntax, semantics, quality, and terms of use of the data offered by the data product. The practitioners proposed implementing data contracts by storing the relevant metadata or defining them as API using standards such as Open API  [^3] https://www.openapis.org/.
Zhamak recommends using the infrastructure as code (IaC) approach to automate the provisioning of the infrastructure of a data product (Dehghani, 2022a). The gray literature also highlighted the importance of IaC (AgileLab, 2021; Limited, 2021; Dehghani, 2021a; Ramos, 2021; Wolf, 2022) and provided several examples of data products that use IaC for deploying them (Nivas Shankar and Hasson, 2021; Orwat, 2021; Raab and Chiapusso, 2021; McKenzie, 2021). IaC employs the source code to manage and provision infrastructure resources and deploy and configure applications, enabling the application of software engineering tools and best practices (Morris, 2020).
Types of Data Products.
@@
Zhamak (Dehghani, 2022b, 2019b, a) mentioned three types of data products: source-aligned, aggregates, and consumer-aligned. The gray literature also uses the same terminology for categorizing data products. Source-aligned data products (Schultze and Wider, 2020b; Menning, 2021; Strenholt and Courtright, 2022; Thoughtworks, 2021; Ramos, 2022; Strengholt, 2022e; Jochen Christ and Harrer, 2022; Strengholt, 2022b) ingest and process data from source systems such as operational databases, external APIs, and sensors. The processed data is then offered as a product to be consumed by the organization’s end-users or downstream data products. The aggregate data products (Dehghani, 2019a; Kumar, 2022; Jochen Christ and Harrer, 2022) compose the data from the upstream data products to produce the value-added data for downstream consumers. Compared to the aggregates, consumer-aligned data products (Schultze and Wider, 2020b; Strenholt and Courtright, 2022; Thoughtworks, 2021; McLarty, 2021; Platter, 2021b; Ramos, 2021; Strengholt, 2022d; Jochen Christ and Harrer, 2022; Strengholt, 2022b) are generally not sold as data products to other domains; instead, they serve data to implement the end-user use cases of the organization.
[…]
@@
An example of data products in different domains, adopted from (Jochen Christ and Harrer, 2022).
Figure 5. An example of data products in different domains, adopted from (Jochen Christ and Harrer, 2022).
@@
Figure 5 shows data products spread across domains. Some data products, e.g., product search, order, and invoice, are source-aligned products. They mostly ingest raw data from the underlying operational data sources (e.g., product and order databases), clean and transform the raw data, and store them locally. They also provide the interfaces to serve their local domain data to other data domains. The aggregate data products, e.g., 360-degree customer views and product recommendations, aggregate data from one or more domains and offer as a value-added product to one or more domains. Some data products, e.g., funnel analytics and CRM dashboards, primarily serve end-users (i.e., consumer-aligned data products).
Characteristics of Data Products.
Zhamak initially introduced six properties for a data product (Dehghani, 2019b) and later added two properties (Dehghani, 2022a). From the gray literature, we tried to identify the practitioner’s perspectives on them.
Discoverable.:
Data consumers should be able to discover the data products in an organization easily. A common approach proposed in the gray literature is to use a central data/product catalog to store the metadata of data products (Moses, 2020; Schultze and Wider, 2020b; Serra, 2021; AgileLab, 2021; JPMorgan, 2021; Pollock, 2021; Nivas Shankar and Hasson, 2021; Vellante, 2021; Wulveryck, 2021; golombek, 2021; Perlov, 2021; Oracle, 2021a; Chinnasamy and Menon, 2021; Keboola, 2021; Rajagopalan, 2021; Das, 2020; Gaur, 2021; Orwat, 2021; Dalvik, 2020; Sequeda, 2021; Nguyen, 2021; Moses, 2021; Shahrin, 2021; Madan, 2021; Raab and Chiapusso, 2021; Oracle, 2021b; Reid, 2021; Wells, 2021; Broda, 2022c, b, a, d; Thoughtworks, 2021; Modi, 2021; Ramos, 2021; Broda, 2022e; Maharana, 2021; AWS, 2021; Velante, 2021; Jochen Christ and Harrer, 2022; Strengholt, 2022c). The data catalog can enable consumers to search for data products using metadata. However, Zhamak (Dehghani, 2022a) does not recommend a central data catalog managed by a separate team, as data products are responsible for deciding and providing their discoverability information. Instead, the data product can offer a discovery interface that can be used to build product search and exploration capabilities at the mesh level.
Interoperable.: